Lane invested $14000 part at 15% and part at 2%. If the total interest at the end of the year is $1060, how much did she invest at 15%. Please help. What formula could I use?
Answer by stanbon(75887)

Lane invested $14000 part at 15% and part at 2%. If the total interest at the end of the year is $1060, how much did she invest at 15%.
------------
Let amount invested at 15% be "x"; interest on this is 0.15x dollars
------------------
Amount invested at 2% is 14000-x: interest on this is 0.02(14000-x)= 280-0.02x
----------------------------------
EQUATION:
interest + interest = 1060
0.15x + (280 -0.02x) = 1060
0.13x = 780
x = $6000 ( amount invested at 15%)
14000-x = 14000-6000 = $8000 (amount invested at 2%)
===============
Cheers,
Stan H.